Our teacher asked us this question today and we were asked to figure it out but no one was able to, so I'm asking someone better than me at this. Here is your problem.
1+2=3
1+2=4
1+2=5
All solutions are true how is this possible?

Answer by ikleyn(52776)

.
Our teacher asked us this question today and we were asked to figure it out but no one was able to,
so I'm asking someone better than me at this. Here is your problem.
All solutions are true how is this possible?
~~~~~~~~~~~~~~~~~~

1+2=3          True
1+2=4          False
1+2=5          False